Multiple Choice questions

Question 2
When line losses are neglected, simpler expressions for line parameters are
obtained. How are the expressions of the exact
1
Expert's answer
2021-08-12T07:27:51-0400

When line losses are neglected (R=G=0) simpler expressions are obtained and certain important

concepts are more easily analyzed. Lossless line Generally the lines are designed to have very small losses. Also for high-frequency studies such as lightning surges, the lossless line approximation will be even accurate, as wL and wC are very large compared to R.

Characteristic impedance Z for a lossless line is derive as below

"Z=\sqrt{\dfrac{R+j\omega{L}}{G+j\omega{C}}}\\
R=G=0\\
\therefore
Z=\sqrt{\dfrac{L}{C}}"